Differential Equation: y^2dx-x(2x+3y)dy=0

1
Expert's answer
2021-11-22T19:22:21-0500
"y^2-x(2x+3y)y'=0"

"y'=\dfrac{y^2}{x(2x+3y)}"

Substitute "y=vx"

"y'=v+v'x"

"v+v'x=\dfrac{v^2x^2}{x(2x+3vx)}"

"v'x=\dfrac{v^2-2v-3v^2}{(2+3v)}"

"\dfrac{2+3v}{2v(v+1)}dv=-\dfrac{dx}{x}"

Integrate


"\int\dfrac{2+3v}{2v(v+1)}dv=-\int\dfrac{dx}{x}"

"\dfrac{2+3v}{v(v+1)}=\dfrac{A}{v}+\dfrac{B}{v+1}=\dfrac{Av+A+Bv}{v(v+1)}"

"A=2, B=1"

"\int\dfrac{2}{v}dv+\int\dfrac{1}{v+1}dv=-\int\dfrac{2dx}{x}"

"2\ln(|v|)+\ln(|v+1|)=-2\ln(|x|)+\ln C"

"v^2(v+1)=\dfrac{C}{x^2}"

"x^2v^2(vx+x)=Cx"

"y^2(x+y)=Cx"